Paths

Version 1.0 February,27 1993

by Enrico Fedrigo

Permission is granted to make and distribute verbatim copies of this manual provided the copyright notice and this permission notice are preserved on all copies.

We don’t guarantee that the programs described in this document are 100% reliable. You are using this material at your own risk. The author can not be made responsible for any damage which is caused by using these programs.

This package is freely distributable, but still copyright by Enrico Fedrigo. This means that you can copy it freely as long as you don’t ask for a more than nominal copying fee. This fee must not be more than US $5.

Permission is granted to include this package in Public-Domain collections, especially in Fred Fish Amiga Disk Library (including CD ROM versions of it). The distribution file may be uploaded to Bulletin Board Systems or FTP servers. If you want to distribute this program you must use the original distribution archive ‘Paths1_0.lha’.

1 Address

Author may be reached at:

Postal Address

                   Enrico Fedrigo
                   via Giacomo Leopardi, 6
                   31033 Castelfranco Veneto (Treviso)
                         ITALY

Electronic Mail (InterNet):

               fedro@sabrina.dei.unipd.it
               fedro@paola.dei.unipd.it

2 Introduction

Paths is a preference editor that fills a lack of WorkBench.

Using Paths you don’t need anymore tens of path lines in your “User-Startup”, it’s all made in a single line (or even less, see Chapter 3 [Installation], page 4).

Paths may be useful both for the boot and for adding or removing global or local paths during your works.

Paths is the second program I made that I decided to leave of public domain, and I’ve been careful following all Commodore standards. Thus the window is an AppWindow font-sensitive, it works with iff files, accepts the command “version”, can be used either from WorkBench or from CLI, and uses the same kind of menus of all the other prefs.

3 Installation

The installation is really easy:

− Copy “Paths” and “Paths.info” in “SYS:prefs”. If your processor is a 68030 or 68040 you may choose the optimized version named “Paths.030”. Copy and rename it in place of normal “Paths”.

− Copy in LOCALE:Catalogs/<language> the file of your language, if is present in the archive. If not, you can fill the empty catalog Paths.ct: if you will send me, I’ll include it in the next version.

− Activate Paths − use option “Read System List” from “Edit” menu − Delete with “Del” button all unmodified default system path assignments in the path list, such

as^{1}:

        C: S: SYS:Tools SYS:Tools/Commodities SYS:Utilities
        SYS:Rexxc SYS:System SYS:Prefs SYS:WbStartup

− Did you add custom paths in your User-Startup or Startup-Sequence?

• If not: − Save your list in “SYS:WbStartup” with “Save As...” from “Project” menu

• If yes: − Save configuration with “SAVE” button − Add before the first call of your custom paths:

               SYS:Prefs/Paths USE

The installation is completed.

1 Warning: you won’t find symbolic devices C:, S: etc., but their full name

4 Use

Paths has either CLI and WB interface.

If you start the program from WB you will be able to change only global paths, while starting it from shell you will be able to change even shell’s own local paths without interfering with global workBench paths. In “Preferences” mode you will modify the path list for the next boot or even the current (local or global) list. You can change mode from ”Preference” option in the “Project” menu.

Important: when changing paths from “Preferences” mode, they will be added (without double paths) to the currents ones; otherwise they will entirely substitute the old ones.

Activating the “EDIT” flag or executing Paths from WorkBench editing window is opened.

Opening the program by icon you can set some tooltypes similar to CLI parameters. Of course these tooltypes may be selected even in the project icons created by Paths and used to load it.

To add a new path just click on ‘ADD’ gadget. You’ll be asked for a new Target. ”DEL” button deletes selected destination.

List order is really important: the directory that includes the common use commands should be first in list. You may change the order simply by selecting one by one the paths and chosing “First”, “Up”, “Down”, “Last”. These actions allow to move the selected path to top of list, up/down of one place and at bottom of list.

Paths allows even to add targets by carrying the path (directory or device) icon above Paths window. This is useful to select several targets simultaneously.

If the icon was a valid preference file, it’ll be used.

By a single click on a TARGET you can modify its path using the gadget string. With a double click you choose it from a file requester.

From menu you can choose other options.

The last three buttons:

•‘SAVE’ Saves configuration on disk (ENVARC:), executes the changes and exits. •‘USE’ Saves configuration on ram (ENV:), executes the changes and exits. The configuration

won’t be available after a reboot.

•‘CANCEL’ Exits without saving or executing assignments.

Obviously, as any preference editor, even Paths may be started once a time: starting a second instance returns only a “popup” of the existent Paths window and its activation.

5 CLI Execution

CLI syntax (available typing ‘Paths ?’):

FROM,EDIT/S,USE/S,SAVE/S,PUBSCREEN/K,BOX3D/S,PREFERENCE/K,GLOBAL/K

−‘FROM’ File to be load instead of “ENV:Sys/Path.prefs”. FROM key may be

omitted.

−‘EDIT’ [Default]

Once loaded prefs file, opens editing window to allow changes

−‘USE’ Once loaded prefs file, executes paths and exits. −‘SAVE’ Once loaded prefs file, saves it on “ENVARC:Sys/Path.prefs”, executes the

paths and exits.

−‘PUBSCREEN’ <pub screen name>

Opens editing window on indicated public screen.

−‘BOX3D’ The gadgets have a contour, there is a NeXT-like box. Selecting Yes or No you choose to use or not it. (This is not an Amiga standard)

−‘PREFERENCE’ By default the program allows to change a preference file and add paths. Selecting this flag you get the exact image of all present system paths (local and global) and you can change, add or remove.

−‘GLOBAL’ By default a shell starting Paths works on local paths. Selecting this flag you can edit global paths.

You may change local/global mode even from the program, changing it with the CYCLE-gadget.

6 ToolTypes

The availables ToolTypes (usable even by Paths’s icons) are:

−‘EDIT’ [Default]

Once loaded prefs file, opens editing window to allow changes

−‘USE’ Once loaded prefs file, executes assignments and exits. −‘SAVE’ Once loaded prefs file, saves it on “ENVARC:Sys/Path.prefs”, executes the

assignments and exits.

−‘CREATEICONS=YES|NO’ [Default = YES] Choose Yes/No whether or not you want to create icons for the files created with ’Save As...’ from menu (see Chapter 7 [Menu], page 9).

−‘PUBSCREEN’ <pub screen name>

Opens editing window on indicated public screen.

−‘BOX3D=YES|NO’ [Default = YES] The gadgets have a contour, there is a NeXT-like box. Selecting Yes or No you choose to use or not it. (This is not an Amiga standard)

You may select EDIT,USE and SAVE even as:

ACTION=EDIT | USE | SAVE

7 Menu

•‘Project’ −‘Open...’ Allows to load a previously saved prefs file. −‘Save As...’

Saves paths-lists. If CREATEICONS is YES, an icon is added to the file. The icon may be used to activate the paths-list. Default tooltype is ‘ACTION=USE’.

−‘Quit’ Quits the program.

•‘Editor’

−‘Read System List’ Reads all present system or custom assigned paths. You lose the currents eventually listed assignments.

−‘Preference’ Switch from Preference mode, in which you can edit preference file and paths will be added to the selected environnement, to the second mode in which you can modify actually present paths in the selected environnement.

−‘Last Saved’ Reads last saved assignments from ENVARC: You lose the currents eventually listed assignments.

−‘Restore’ Loads last valid assignments from ENV: You lose the currents eventually listed assignments.

•‘Settings’

−‘Create Icons?’ (Flag)

Tells whether or not to add an icon to the file generated by Paths with “Save As...” option. When starting takes tooltype (see Chapter 6 [ToolTypes], page 8) CREATEICONS value. Default is Yes.

8 Distribution

The complete distributed Paths package includes:

Paths1.0.info
Paths1.0/Paths  • 680x0 version
Paths1.0/Paths.030  • 68030/881 version
Paths1.0/Paths.info
Paths1.0/Docs.info
Paths1.0/Docs/Italiano.doc
Paths1.0/Docs/Italiano.doc.info
Paths1.0/Docs/Italiano.guide
Paths1.0/Docs/Italiano.guide.info
Paths1.0/Docs/Italiano.dvi
Paths1.0/Docs/Italiano.dvi.info
Paths1.0/Docs/Italiano.texi
Paths1.0/Docs/Italiano.texi.info
Paths1.0/Docs/English.doc  • plain ASCII docs
Paths1.0/Docs/English.doc.info
Paths1.0/Docs/English.guide  • AmigaGuide docs
Paths1.0/Docs/English.guide.info
Paths1.0/Docs/English.dvi  • DVI docs
Paths1.0/Docs/English.dvi.info
Paths1.0/Docs/English.texi  • TeXinfo sources
Paths1.0/Docs/English.texi.info
Paths1.0/Locale/Catalogs/Italiano/Paths.catalog
Paths1.0/Locale/Catalogs/Paths.ct

These files MUST be present.

Chapter 9: Future develops 11

9 Future develops

Now the program works fine and is reliable, next version will be optimized in memory usage. I intend to write the next version of Paths in C++ and I’m beginning to study some objects for

programming in AmigaDOS.

Chapter 10: Versions 12

10 Versions

Paths1.0: first public release.

11 Greetings

My greetings for the help (tests and suggestions) are for:^{1}:

Alessandro Borille    paride@sabrina.dei.unipd.it
Eugenio Castellani    aj2@sabrina.dei.unipd.it
Fabio D'Alessio       phobia@sabrina.dei.unipd.it
Sergio Fantinel       ik3fan@paola.dei.unipd.it
Jacopo Farinato       redi@sabrina.dei.unipd.it
Gianluigi Nusca       gigio@paola.dei.unipd.it
Emanuele Verza        alien@sabrina.dei.unipd.it

Specially Alessandro and Eugenio for italian to english translation.

1 alphabetically listed